OpenStudy (wintersuntime):

Rewrite this equation in y= mx+b -18x+3y= -12

OpenStudy (wintersuntime):

@PillowFace750

OpenStudy (pillowface750):

@JacoptimalRBLX @pink33 good luck

OpenStudy (anonymous):

The slope-intercept form is y=mx+b, where m is the slope and b is the y-intercept. y=mx+b

OpenStudy (anonymous):

y=−2/3x+2

OpenStudy (wintersuntime):

thanks

OpenStudy (anonymous):

Np.

rebeccaxhawaii (rebeccaxhawaii):

solve for y